# Green Operations of SWIPT Networks: The Role of End-User Devices

### v1

@@ Abstract
Internet of Things (IoT) devices often come with batteries of limited capacity that are not easily replaceable or rechargeable, and that constrain significantly the sensing, computing, and communication tasks that they can perform. The Simultaneous Wireless Information and Power Transfer (SWIPT) paradigm addresses this issue by delivering power wirelessly to energy-harvesting IoT devices with the same signal used for information transfer. For their peculiarity, these networks require specific energy-efficient planning and management approaches. However, to date, it is not clear what are the most effective strategies for managing a SWIPT network for energy efficiency. In this paper, we address this issue by developing an analytical model based on stochastic geometry, accounting for the statistics of user-perceived performance and base station scheduling. We formulate an optimization problem for deriving the energy optimal configuration as a function of the main system parameters, and we propose a genetic algorithm approach to solve it. Our results enable a first-order evaluation of the most effective strategies for energy-efficient provisioning of power and communications in a SWIPT network. We show that the service capacity brought about by users brings energy-efficient dynamic network provisioning strategies that radically differ from those of networks with no wireless power transfer.
@@ I Introduction
The advent of the Internet of Things (IoT) paradigm has significantly impacted the energy consumption of Radio Access Networks (RAN) [2]. On one side, the proliferation of IoT devices is leading to an exponential increase in the number of connected devices and in the amount of data generated and exchanged [3]. On the other, the wide heterogeneity of IoT communication requirements and traffic patterns, together with the irregularity in the profiles of device activity and sleep cycles, induce patterns of energy consumption within the RAN that differ substantially from those of traditional, broadband-centric cellular networks [4, 5], with a strong impact on the way the network is planned and operated.
@@
This impact is further amplified when the cellular network does not only deliver connectivity to all its users but also power to IoT devices. In this work, we consider scenarios in which IoT devices harvest RF energy from the environment (passive energy harvesting - EH) as well as from the signal transmitted by their serving base station (active EH) [6]. Such configurations are interesting as they allow for overcoming the power budget constraints that limit the potential of battery-operated IoT in many application domains. Indeed, these limitations often imply tight constraints on the amount of sensing, computing, and actuation that IoT devices can perform, and negatively affect their availability.
@@
Among the available technologies for wireless power delivery to EH devices, simultaneous wireless information and power transfer (SWIPT) [7, 8, 9, 10, 11] is of particular interest, as it exploits the same cellular infrastructure used to deliver data. This feature makes it easier and more cost-effective to deploy than alternative approaches. However, different from traditional single-service cellular networks (i.e., networks with no wireless power transfer), energy efficiency in SWIPT networks encompasses both the energy consumed by BSs and that consumed by IoT users, which, for their ubiquity and rising numbers, are poised to play a key role in determining the overall energy footprint of the network. Thus, SWIPT networks require the development of specific approaches for energy-efficient network design and management, which may substantially differ from traditional, single-service, RANs.
Several works on energy efficiency in SWIPT [12, 13, 14] consider simple scenarios composed of a single base station (BS). They focus on accounting for the impact of the interplay between power and information delivery on the energy footprint of the network. Other works ([15, 16]) consider specific BS deployment scenarios serving a population of EH IoT devices, and they propose some heuristics for saving energy in those configurations. However, these algorithm and scenario-specific approaches do not allow drawing general conclusions on the impact of the main system parameters on energy efficiency in a generic SWIPT network, and thus getting an idea of the actual potential for energy saving of dynamic network management in SWIPT. Indeed, these aspects are key for supporting any decision on how to effectively plan and manage such networks.
Another set of results focuses on a system-level analysis of SWIPT networks based on the tools of stochastic geometry [17, 18, 19]. Crucially, however, none of these works account for the effects of resource scheduling among users, nor do they address the key issue of determining the energy-optimal strategies for SWIPT network provisioning and QoS-aware dynamic management.
@@
In this paper, we address these issues and propose an analytical framework based on stochastic geometry (SG) for system-level performance evaluation of SWIPT cellular networks, which accounts for resource scheduling among broadband and IoT EH users. The use of SG allows focusing on the average behavior of the system over many realizations of the process of UE (user equipment) and BS spatial distributions. This allows characterizing the potential for QoS-aware energy optimization of SWIPT-enabled networks and the main trade-offs between user-perceived performance and resource utilization in a scenario-independent manner. Specifically, the results of this paper are as follows:
  • We derive analytical expressions for the key statistics of the main performance indicators of a SWIPT network.
  • We formulate the problem of deriving a configuration (in terms of density of active BSs, split ratio, and transmit power) that minimizes the mean total power consumption of the network while achieving a target QoS level for both information transfer (in downlink and uplink) and wireless power delivery.
@@
  • We propose a genetic algorithm (GA) approach for deriving energy-efficient SWIPT network configurations, achieving the target user-perceived performance levels.
@@
  • We characterize the key performance and resource efficiency trade-offs of the configurations resulting from our heuristic. This allows observing some rather surprising effects, such as, in some cases, a decrease in the optimal density of BSs required to serve an increasingly dense population of UEs and the irrelevance of active EH for very dense populations of IoT devices.
The paper is organized as follows. In Section II, we introduce the related works on system modeling for SWIPT networks. In Section III, we present the system model, and in Section IV, we derive a model for user-perceived performance for information and power transfer. In Section V, we formulate an optimization problem for the maximization of the energy efficiency of the network. In Section VI we assess numerically our results, and in Section VII, we conclude the paper with some final remarks.
@@ II Related works
The growing interest of the wireless communication community towards the SWIPT paradigm is due to its potential for enabling ubiquitous and self-sustained cellular-based IoT [7, 8, 9, 10, 11]. Moreover, differently from IoT energy harvesting techniques based on solar or wind, those based on active wireless power transfer have the advantage of being stable and available at any time [6]. Several technical challenges however still need to be addressed for the practical viability of SWIPT networks, related to energy-harvesting transceiver and algorithm design, system integration, protocol design, and energy-efficient network planning and operations [20, 21].
@@
Initial works on energy efficiency in SWIPT networks considered scenarios composed of a BS or a broadband unit with several remote radio heads, and a population of EH devices. These works aim to optimize the system configuration in terms of energy efficiency [12, 13, 22]. Among these, [12] characterizes the data rate vs. energy trade-off in a single SWIPT BS scenario. [13] proposes an approach for optimizing transmit power allocation to each user. [22] proposes a Markov chain model to derive the joint power and connectivity outage probability. [23] investigates strategies for transmit-power-efficient resource allocation. However, the simple scenarios considered in these works do not allow for extending the results or extrapolating them to different and more realistic configurations, nor do they account for the impact of the statistics of BS and user distribution.
@@
Algorithms for deriving energy-optimal configurations of sample SWIPT systems have been proposed for networks with CoMP [24], in NOMA networks [25], in full duplex settings [26], as well as in scenarios with two-way relays [27]. Among these, [15] proposes an algorithm to minimize the total energy consumption of a set of BSs and a population of IoT devices, over sub-carrier and transmit power allocation. However, the strategies proposed in these works focus on deriving the optimal configuration for a specific set of deployed BS, with their specific layout and user spatial distribution. As such, they do not allow evaluating the overall potential for energy-saving enabled by their schemes in a scenario-independent manner. Thus, they leave open the more general issue of what are the most effective management strategies to achieve system-level energy proportionality in a SWIPT network, as a function of the main system parameters.
@@
Similarly to ours, another set of works ([17, 18, 28]) performs system-level analysis, modeling the average behavior of a SWIPT network, such as the data rate vs. energy trade-off (e.g., see [18]), or optimizing the D2D successful transmission rate [28]. These works however do not account for the effects of resource allocation and scheduling among users, such as the statistics of the sharing of BS time across all associated users, which are key for an accurate modeling of user-perceived performance. Thus, they do not enable an accurate and realistic characterization of the main performance trade-offs, and in particular of the strategies for QoS-aware energy-efficient operation of SWIPT networks.
@@ III System Model
We assume that BSs are distributed in space according to a homogeneous planar Poisson Point Process (PPP) with a density equal to $\lambda_{b}$ BSs per $km^{2}$. UEs are distributed in space according to a homogeneous PPP with an intensity equal to $\lambda_{u}$ UEs per $km^{2}$. UEs are either broadband (BB) terminals, or IoT (Internet of Things) devices. We assume the latter are a fraction $\gamma$ of the total number of UEs.
@@
  • static power splitting (SPS), by which a fraction $\mathbf{\nu}$ (the power split ratio, with $0\leq\nu\leq 1$) of the power received at any time instant by an IoT device is harvested.
  • dynamic power splitting (DPS), by which a fraction $\mathbf{\nu}$ of the power transmitted by any source during the time the serving base station devotes to a given IoT device is harvested. During the rest of the time, all of the received power is harvested at the node.
@@
  • a linear model, by which $h=\xi h_{in}$, with $\xi\leq 1$; and
@@
$\bar{\tau}_{j}$, $j=d,u$
Mean ideal per-bit delay in downlink/uplink ($s^{-1}$)
$\tau_{j}$, $j=d,u$
Per-bit delay in downlink/uplink ($s^{-1}$)
$\tau_{j}^{0}$, $j=d,u$
Target per-bit delay in downlink/uplink ($s^{-1}$)
$\xi$
EH conversion efficiency ratio (linear model)
$h_{0}$
Minimum harvested power required by each IoT device ($W$)
$\mu$
Maximum acceptable ratio of IoT users which harvest less than $h_{0}$ W
$\phi$
Fraction of time in which IoT devices are active
$\lambda_{b}$
Mean BS density ($m^{-2}$)
$\lambda_{u}$
Mean density of user terminals ($m^{-2}$)
$\gamma$
Fraction of user terminals which are IoT devices
$\eta$
Fraction of the time dedicated by a BS to serve an IoT device in downlink used for harvesting energy from the signal received from the base station
$\nu$
Fraction of the power received at any time instant by an IoT device which is used for energy harvesting
$\beta_{j}$, $j=d,u$
GPS weight of the amount of time spent by a BS in idle mode in downlink/uplink
$w_{d}$
Ratio between the amount of time dedicated to serving a BB UE and the one dedicated to serving an IoT user in downlink
$\delta_{j}$, $j=d,u$
Ratio between the BS time dedicated to a BB user, and the base station time dedicated to an IoT user for data transfer in downlink/uplink
$k$
Frequency reuse factor
$\alpha$
Path loss exponent
$P$
BS transmit power ($W$)
$G$
BS antenna gain
$L$
BS antenna loss (out of beam)
@@
  • Information transfer: For all UEs, the end-user performance metric is the per-bit delay $\tau$ of data transfers, defined as the inverse of the short-term user throughput, i.e., the actual rate at which the user is served.
@@ III-A Channel and Service Model
Our channel model only takes into account distance-dependent path loss. The incorporation of the effects of fading and shadowing would not alter our approach, and it is left for future consideration. We assume that random frequency reuse is in place, with reuse factor $k$. That is, every BS is assigned one out of $k$ frequency bands with equal probability.
We assume that UEs are associated with the BS that provides the largest SINR at the user location. We consider urban scenarios, where the high capacity demand justifies the use of strategies for energy efficient network planning and management, and where the assumption of high attenuation (with exponent $\alpha\geq 3$) typically holds. In these settings, as no fading is considered and all BSs use the same transmit power, assuming that users associate to the closest BS is a reasonable approximation [19].
@@
As users are assumed to be distributed uniformly in space, and as base station time is supposed to be shared in equal parts across users of the same type, the mean power received by a user while it is not being served from the base station to which it is associated (located at a distance $r$) is $PL_{g}r^{-\alpha}$, with $L_{g}=\frac{Ga+L(360-a)}{360}$. Thus $L_{g}r^{-\alpha}$ is the mean attenuation with which the power transmitted by the serving base station is received by a user when it is not being served.
@@
$C(r,P,G,I)=(B/k)\log_{2}\left(1+\frac{PGr^{-\alpha}}{N_{0}+I(r,k)}\right)$ (2)

### v3

@@ Abstract
Internet of Things (IoT) devices often come with batteries of limited capacity that are not easily replaceable or rechargeable, and that constrain significantly the sensing, computing, and communication tasks that devices can perform. The Simultaneous Wireless Information and Power Transfer (SWIPT) paradigm addresses this issue by delivering power wirelessly to energy-harvesting IoT devices with the same signal used for information transfer. For their peculiarity, these networks require specific energy-efficient planning and management approaches. However, to date, it is not clear what are the most effective strategies for managing a SWIPT network for energy efficiency. In this paper, we address this issue by developing an analytical model based on stochastic geometry, accounting for the statistics of user-perceived performance and base station scheduling. We formulate an optimization problem for deriving the energy-optimal configuration as a function of the main system parameters, and we propose a genetic algorithm approach to solve it. Our results enable a first-order evaluation of the most effective strategies for energy-efficient provisioning of power and communications in a SWIPT network. We show that the service capacity brought about by users brings energy-efficient dynamic network provisioning strategies that radically differ from those of networks with no wireless power transfer.
@@ I Introduction
The advent of the Internet of Things (IoT) paradigm has significantly impacted the energy consumption of Radio Access Networks (RAN) [2]. On the one side, the proliferation of IoT devices is leading to an exponential increase in the number of connected devices and in the amount of data generated and exchanged [3]. On the other, the wide heterogeneity of IoT communication requirements and traffic patterns, together with the irregularity in the profiles of device activity and sleep cycles, induce patterns of energy consumption within the RAN that differ substantially from those of traditional, broadband-centric cellular networks [4, 5], with a strong impact on the way the network is planned and operated.
@@
This impact is further amplified when the cellular network does not only deliver connectivity to all its users but also power to IoT devices. In this work, we consider scenarios in which IoT devices harvest RF energy from the environment (passive energy harvesting - EH) as well as from the signal transmitted by their serving base station (active EH) [6]. Such configurations are interesting as they allow for overcoming the power budget constraints that limit the potential of battery-operated IoT devices in many application domains. Indeed, these limitations often imply tight constraints on the amount of sensing, computing, and actuation that IoT devices can perform, negatively affecting their availability.
@@
Among the available technologies for wireless power delivery to EH devices, simultaneous wireless information and power transfer (SWIPT) [7, 8, 9, 10, 11] is of particular interest, as it exploits the same cellular infrastructure used to deliver data. This feature makes it easier and more cost-effective to deploy than alternative wireless power delivery approaches. However, differently than traditional single-service cellular networks with no active power transfer, managing SWIPT networks for energy efficiency implies accounting for the interdependence between the energy consumed by BSs and that consumed by IoT users which, for their ubiquity and rising numbers, are poised to play a key role in determining the overall energy footprint of the network. Thus, SWIPT networks require the development of specific approaches for energy-efficient network design and management, which may substantially differ from traditional, single-service RANs.
The issue of energy efficiency in SWIPT is being actively investigated [12, 13, 14]. However, these works consider simple network scenarios, composed of a single base station (BS). They focus on accounting for the impact of the interplay between power and information delivery on the energy footprint of the BS. Other works ([15, 16]) consider a set of BSs deployed according to a specific layout, and they propose heuristics for saving energy in those configurations. Being tied to the specifics of the scenario considered, these heuristic approaches do not allow drawing general conclusions on the main features and limitations of energy-optimal management strategies in SWIPT networks. Another line of research focuses on applying the tools of stochastic geometry (SG) [17, 18, 19] for modeling system-level average performance in SWIPT networks. SG allows focusing on the average behavior of the system over many realizations of the process of UE (user equipment) and BS spatial distributions, making it possible to analyze the trade-offs between different potential objectives such as maximizing coverage, maximizing throughput, and maximizing energy harvesting, in ways that would otherwise be very difficult or way less accurate. Crucially, however, none of these results account for the effects of resource scheduling among users on user-perceived performance, nor do they address the key issue of characterizing energy-optimal strategies for QoS-aware (for both power and information delivery) dynamic network provisioning, as a function of traffic and energy demand. Thus, it is still unclear what is the actual potential for energy efficiency of strategies which dynamically tune the configuration of SWIPT networks, and what is the impact of optimizing the main system parameters on the energy consumed by IoT users and BSs.
@@
In this paper, we address these issues and propose a stochastic geometry analytical framework for modeling the relationship between energy consumption, user perceived performance, and the main system parameters in SWIPT cellular networks, which accounts for QoS-aware resource scheduling among broadband and IoT EH users. Our framework allows characterizing in a scenario-independent manner the potential for QoS-aware energy efficiency in these networks, as well as the main trade-offs between user-perceived performance in the communication and power delivery services, resource utilization, and overall energy consumption of the network. Specifically, the contributions of this paper are as follows:
  • We derive a set of analytical results for the key statistics of the main performance indicators of a SWIPT network, which account for resource scheduling among broadband and IoT EH users, and allow modelling the impact of the main system parameters on network energy consumption;
  • We formulate an optimization problem to determine the potential energy savings achievable by tuning the main network parameters while guaranteeing a target user-perceived performance for both information transfer (in downlink and uplink) and wireless power delivery.
@@
  • We elaborate a Genetic Algorithm (GA) approach to derive energy-efficient SWIPT network configurations which achieve the target QoS levels;
@@
  • We validate numerically our approach, shedding light on several aspects of the tradeoff between user-perceived performance and resource efficiency in a SWIPT network. Our analysis allows observing some unexpected and rather surprising effects such as, in some cases, a decrease in the optimal density of BSs required to serve an increasingly dense population of UEs, and the irrelevance of active power delivery for very dense populations of IoT devices.
The rest of this paper is organized as follows. Section II reviews the relevant state of the art. In Section III, we present the system model, and in Section IV our analytical approach for user-perceived performance modelling for both information and power transfer. In Section V, we present our formulation of the optimization problem and the GA heuristic for solving it efficiently. In Section VI we assess numerically our results. Finally, Section VII concludes the paper.
@@ II Related works
The growing interest towards the integration of wireless power transfer in cellular networks is due to its potential for enabling truly ubiquitous and self-sustained cellular-based IoT [7, 8, 9, 10, 11]. Indeed, differently from other energy harvesting techniques (e.g., based on solar, or solely on passive RF harvesting) those based on active wireless power transfer have the advantage of being stable and available at any time [6]. Several technical challenges however still need to be addressed for the practical viability of SWIPT networks, related to energy-harvesting transceiver and algorithm design, system integration, protocol design, and energy-efficient network planning and operations [20, 21].
@@
Initial works on energy efficiency in SWIPT networks considered scenarios composed of a BS or a broadband unit with several remote radio heads, and a population of EH devices. These works aim to optimize the system configuration in terms of energy efficiency [12, 13, 22]. Among these, [12] characterizes the data rate vs. energy trade-off in a single SWIPT BS scenario. For the same setting, [13] proposes an approach for optimizing transmit power allocation to each user. [22] elaborates a Markov chain model to derive the joint power and connectivity outage probability. [23] investigates strategies for transmit-power-efficient resource allocation. All of these works provide a first insight into the basic performance patterns of a SWIPT system. However, they are based on single BS configurations, which are not representative of the average performance of a whole SWIPT network, and they do not account for the impact of the statistics of BS and user distribution. In particular, they lack accurate modelling of the effects of co-channel interference and passive RF power transfer from BSs other than the serving one, as well as from users associated with them.
@@
Another set of results focuses on sample SWIPT network configurations, with a given set of BS and a given spatial layout of BS and of users. They propose algorithms for energy-optimal configurations, for networks with CoMP [24], in NOMA networks [25], in full duplex settings [26], as well as in scenarios with two-way relays [27]. Among these, [15] proposes an algorithm to minimize the total energy consumption of a set of BSs and a population of IoT devices, over sub-carrier and transmit power allocation. However, the strategies proposed in these works focus on deriving the optimal configuration for a sample BS deployment, with a specific layout and with a given user spatial distribution. Being tied to a specific setup, they do not allow drawing general considerations on the overall potential of these schemes for energy-saving.
@@
Recently, stochastic geometry (SG) [17, 18, 19, 28] has emerged as an effective modelling approach for a stochastic characterization of the performance patterns in a wireless network. SG allows focusing on the average behaviour of the system over many realizations of the process of UE (user equipment) and BS spatial distributions. In SWIPT networks, SG has been used to characterize several performance aspects, such as the data rate vs. energy trade-off [18], or the optimization of the D2D successful transmission rate [28], to name a few. These works however, often for the sake of analytical tractability, do not account for the effects of resource allocation and scheduling among users (such as the statistics of the sharing of BS time across all associated users) which are key for accurate stochastic modelling of the performance perceived by the user. As such, they do not enable an accurate and realistic characterization of the main trade-offs between the energy consumed by the network, user-perceived performance (for both power and information delivery), and resource utilization. Thus, they leave open the more general issue of determining the most effective QoS-aware dimensioning and tuning strategies to achieve system-level energy proportionality in a SWIPT network, as a function of the main system parameters.
In the present paper, we address these issues by proposing an SG modelling framework which captures the relationship between energy consumption, user-perceived performance, and the main system parameters in SWIPT cellular networks. Our framework allows characterizing in a scenario-independent manner the potential for QoS-aware energy efficiency in these networks, as well as the main trade-offs between user-perceived performance in the communication and power delivery services, resource utilization, and overall energy consumption of the network. Our analytical approach overcomes the limitations of existing results, as follows:
  • It allows an estimation of the potential energy savings achievable in a SWIPT network through QoS-aware (for both energy and data transfer) tuning of network parameters.
  • It accounts for the effects of base station resource allocation and scheduling among users on the statistics of user-perceived QoS.
In doing so, our work allows a first characterization of the performance-energy tradeoff, revealing some new and rather unexpected patterns related to the impact of the population of user devices on the overall energy efficiency of the network.
@@ III System Model
We utilize stochastic geometry to model the spatial distribution of network elements in our system. Specifically, we assume that BSs are distributed in space according to a homogeneous planar Poisson Point Process (PPP) with a density equal to $\lambda_{b}$ BSs per $km^{2}$. UEs are distributed in space according to a homogeneous PPP with an intensity equal to $\lambda_{u}$ UEs per $km^{2}$. UEs are either broadband (BB) terminals, or IoT (Internet of Things) devices. We assume the latter are a fraction $\gamma$ of the total number of UEs.
Figure 1: Outline of the system model for the SWIPT wireless network considered in this work.
@@
  • Static Power Splitting (SPS). In this operating mode, the receiver antenna at every IoT device is followed by a splitter, which sends a fraction $\mathbf{\nu}$ (the power split ratio, with $0\leq\nu\leq 1$) of the total signal power received at any time instant to the RF harvesting electronics. The rest of the received signal power is instead used for decoding information.
  • Dynamic Power Splitting (DPS). In this operating mode, the receiver architecture is the same as in the SPS mode. The only difference is that, in DPS, at every IoT user the power split ratio is $\mathbf{\nu}$ when the serving base station is transmitting to that user, and it is equal to one for the rest of the time. That is, when an IoT node is not receiving data from its serving base station, all the received power (from any RF source, including all BSs and users) is fed to the RF harvesting electronics.
@@
  • a linear model, by which $\Theta(h_{in})=\xi h_{in}$, with $\xi\leq 1$; and
@@
$\bar{\tau}_{j}$, $j=d,u$ Mean ideal per-bit delay in downlink/uplink ($s^{-1}$)
$\tau_{j}$, $j=d,u$ Per-bit delay in downlink/uplink ($s^{-1}$)
$\tau_{j}^{0}$, $j=d,u$ Target per-bit delay in downlink/uplink ($s^{-1}$)
$\xi$ EH conversion efficiency ratio (linear model)
$h_{0}$ Minimum harvested power required by each IoT device ($W$)
$\mu$ Maximum acceptable ratio of IoT users which harvest less than $h_{0}$ W
$\phi$ Fraction of time in which IoT devices are active
$\lambda_{b}$ Mean BS density ($m^{-2}$)
$\lambda_{u}$ Mean density of user terminals ($m^{-2}$)
$\gamma$ Fraction of user terminals which are IoT devices
$\eta$ Fraction of the time dedicated by a BS to serve an IoT device in downlink used for harvesting energy from the signal received from the base station
$\nu$ Fraction of the power received at any time instant by an IoT device which is used for energy harvesting
$\beta_{j}$, $j=d,u$ GPS weight of the amount of time spent by a BS in idle mode in downlink/uplink
$w_{d}$ Ratio between the amount of time dedicated to serving a BB UE and the one dedicated to serving an IoT user in downlink
$\delta_{j}$, $j=d,u$ Ratio between the BS time dedicated to a BB user, and the base station time dedicated to an IoT user for data transfer in downlink/uplink
$k$ Frequency reuse factor
$\alpha$ Path loss exponent
$P$ BS transmit power ($W$)
$G$ BS antenna gain
$L$ BS antenna loss (out of beam)
@@
  • Information transfer: For all UEs, the end-user performance metric is the per-bit delay $\tau$ of data transfers. It is the time required to transfer a single bit, and it is thus the inverse of the short-term throughput, i.e., of the rate at which data is transferred.
@@ III-A Channel and Service Model
Our channel model only takes into account distance-dependent path loss. Incorporating the effects of fading and shadowing would not alter our approach, and is left for future consideration. We assume that random frequency reuse is in place, with reuse factor $k$. That is, every BS is assigned one out of $k$ frequency bands with equal probability.
We assume that UEs are associated with the BS that provides the largest SINR at the user location. We consider urban scenarios, where the high capacity demand justifies strategies for energy-efficient network planning and management, and where the assumption of high attenuation (with exponent $\alpha\geq 3$) typically holds. In these settings, as no fading is considered and all BSs use the same transmit power, assuming that users associate to the closest BS is a reasonable approximation [19].
@@
Since users are assumed to be distributed uniformly in space, and base station time is supposed to be shared in equal parts across users of the same type, the mean power received by a user while it is not being served from the base station to which it is associated (located at a distance $r$) is $PL_{g}r^{-\alpha}$, with $L_{g}=\frac{Ga+L(360-a)}{360}$. Thus $L_{g}r^{-\alpha}$ is the mean attenuation with which the power transmitted by the serving base station is received by a user when it is not being served.
@@
$C(r,P,G,I)=\frac{B}{k}\log_{2}\left(1+\frac{PGr^{-\alpha}}{N_{0}+I(r,k)}\right)$ (2)